can anyone solve these for me?
1) 5^(x+2)=8.5
2) 6.5^x = 8.4
i need to know what X is in both of these and the method for figuring them out please
thanks

Answer by lwsshak3(6761)

Whenever you have the variable in the exponent, logs are needed to solve.
1) 5^(x+2)=8.5
(x+2)log(5)=log(8.5)
(x+2)=log(8.5)/log(5)≈1.3297
x=1.3297-2≈-.6703
..
2) 6.5^x = 8.4
x=log(8.4)/log(6.5)
x≈1.137